OpenStudy (anonymous):

I need help finding the indefinite integral (1 to 2) of 1/X^2 (dx).

OpenStudy (anonymous):

hey im not sure how i should use the quotient rule with anti derivs.

OpenStudy (jamesj):

So you want to find the anti-derivative of f(x) = 1/x^2. Call it F(x), where F'(x) = f(x) = 1/x^2

OpenStudy (jamesj):

Once you've found F(x), then your integral is \[ \int_1^2 \frac{1}{x^2} dx = F(2) - F(1) \]

OpenStudy (jamesj):

In other words, what function when differentiated gives 1/x^2 ?

OpenStudy (anonymous):

oh i c mental block how do i find that

OpenStudy (jamesj):

hint: what's the derivative of 1/x ?

OpenStudy (anonymous):

derivative? why not antiderivative? (sorry i need help with this stuff too)

OpenStudy (anonymous):

it is -1/x^2?

OpenStudy (jamesj):

We are looking for the function F(x) with the property that \[ F\ '(x) = 1/x^2. \] Then by the Fundamental Theorem of Calculus \[ \int_1^2 \frac{1}{x^2} dx = F(2) - F(1) \] Now another way to write this \[ F \ '(x) = x^{-2} \] Let's take a step back. What is the derivative of \[ g(x) = x^n \]

OpenStudy (anonymous):

its nx^e?

OpenStudy (anonymous):

nx^9n-1)

OpenStudy (anonymous):

ya what she said

OpenStudy (anonymous):

sorry nx^(n-1)

OpenStudy (anonymous):

its (n+1)x^n+1/n+1 if that makes sense lol

OpenStudy (anonymous):

(1/(n+1))x^(n+1)

OpenStudy (jamesj):

Yes, if g(x) = x^n, then \[ g′(x)=nx^{n−1} \]. Hence, what is the anti-derivative of h(x)=xn I.e., what is the function H such that H' = x^n ?

OpenStudy (anonymous):

yeah that one lol

OpenStudy (jamesj):

Right ... the antiderivative of g(x) = x^n is \[ G(x) = \frac{1}{n+1} x^{n+1} \] Given that, what is the antiderivative of \[ f(x) = 1/x^2 = x^{-2} \]

OpenStudy (anonymous):

x^1?

OpenStudy (anonymous):

oh!! now i get it!!!

OpenStudy (anonymous):

oh sorry -1x ?

OpenStudy (anonymous):

thank you thank you!

OpenStudy (jamesj):

No, f(x) = x^-2. So n = -2. Use now the formula above to find the anti-derivative F(x)

OpenStudy (anonymous):

hmm -1x^-1???

OpenStudy (jamesj):

Right. \[ F(x) = \frac{1}{n+1} x^{n+1} = \frac{1}{-2+1} x^{-2+1} = (-1) x^{-1} = -1/x \]

OpenStudy (jamesj):

Therefore \[ \int_1^2 \frac{dx}{x^2} \] equals what?

OpenStudy (anonymous):

lol hold on sec

OpenStudy (jamesj):

\[ \int_1^2 \frac{dx}{x^2} = F(2) - F(1) \] where F'(x) = 1/x^2

OpenStudy (anonymous):

you explain soo well just new for me trying to get it visualy

OpenStudy (jamesj):

Sketch the graph of f(x) = 1/x^2. We're calculating the tea under the curve between x=1 and x=2.

OpenStudy (jamesj):

So whatever else the answer is, as the area is above the x-axis, the area must be a positive number.

OpenStudy (anonymous):

ahh i c

OpenStudy (anonymous):

so this -1(2)^-1 - -1(1)^-1

OpenStudy (jamesj):

ok. Simplify that.

OpenStudy (jamesj):

you'll find it easier if you think of F(x) as F(x) = -1/x.

OpenStudy (anonymous):

oh i c

OpenStudy (anonymous):

1/2!?

OpenStudy (jamesj):

Yes F(2) - F(1) = -1/2 - (-1/1) = -1/2 + 1 = 1/2

OpenStudy (anonymous):

yessssssssss finally thanks soo much

OpenStudy (jamesj):

The derivative of 1/x is so important, you should learnt it cold. The antiderivative of 1/x^2 is almost as important and you should learn that as well.

OpenStudy (anonymous):

will do :)

OpenStudy (jamesj):

For the record, what is the anti-derivative of -1/x^2 ?

OpenStudy (anonymous):

is it x^2? the other was negative

OpenStudy (jamesj):

1/x^2 = x^-2

OpenStudy (anonymous):

no wait its possitive so -1/2?

OpenStudy (jamesj):

what's the anti-derivative of 1/x^2 ? It's what we just used for your problem.

OpenStudy (anonymous):

its x^-2

OpenStudy (anonymous):

sorry confused it was -1X^-1

OpenStudy (jamesj):

1/x^2 = x^-2 But what is its antiderivative? I.e., what is the function F such that F'(x) = 1/x^2 ?

OpenStudy (anonymous):

was it this -1x^-1?

OpenStudy (jamesj):

It is -1/x, yes. That is, if F(x) = -1/x, then F'(x) = 1/x^2. Ok. Now, what is the antiderivative of f(x) = -1/x^2 ?

OpenStudy (anonymous):

its 1/x?lol i know i have fail memory

OpenStudy (jamesj):

Yes. If F(x) = 1/x, then F'(x) = -1/x^2. ok. ttyl

OpenStudy (anonymous):

Brilliant much thankks for giving me time man